OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 1.26x + 2.16 = 7.71 A. 4.3 B. 9.6 C. 4.5 D. 4.4

OpenStudy (lyrae):

1. Subtract 2.16 from both sides 2. Divide both sides by 1.26

OpenStudy (anonymous):

?

OpenStudy (anonymous):

d

OpenStudy (lyrae):

1.26x + 2.16 = 7.71 Step 1 1.26x + 2.16 - 2.16 = 7.71 - 2.16 1.26x = 5.55 Step 2 1.26x/1.26 = 5.55/1.26 x = 4.40476190476. . . Round to nearest tenth x = 4.4

OpenStudy (lyrae):

Yes :)

OpenStudy (anonymous):

cool thanks

OpenStudy (anonymous):

:)